OpenStudy (partha):

is it possible 1=2 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the value one is never equal to the value two.

OpenStudy (anonymous):

there is a 'proof 'but it is flawed

OpenStudy (anonymous):

using faulty logic yes ^^ however using sound mathematical proofs it is impossible to show that.

OpenStudy (partha):

how?

OpenStudy (anonymous):

but .9999(rep) does equal 1 as you use calculus

OpenStudy (partha):

yes that's correct...

OpenStudy (anonymous):

i dont think the purpose of this site is to teach incorrect math using division by 0 to show 1=2 when it doesnt ; )

OpenStudy (anonymous):

ok try this - let x =1 then x-1 = 2(x-1) now divide both sides by x-1:- 1 = 2

OpenStudy (partha):

from every incorrect there's is a correct ans dear...so for that...

OpenStudy (anonymous):

http://uncyclopedia.wikia.com/wiki/1%3D2#Proof

OpenStudy (anonymous):

For jimmy's answer above, the result is wrong because he's dividing by 0.

OpenStudy (partha):

@dhatraditya ..thanks for the link